OpenStudy (anonymous):

(medal given) Which of the following is not a Pythagorean triple? A. 28, 45, 53 B. 16, 63, 65 C. 13, 84, 85 D. 11, 61, 62

OpenStudy (mathstudent55):

For each choice, square each of the three numbers. Then add the two smaller squares. If the sum of the two smaller squares equals the third square, it's a Pythagorean triple.

OpenStudy (agent0smith):

Do you know how to find out? Use the Pythagorean theorem \[\Large a^2 + b^2 = c^2\]and check if it's true.

OpenStudy (mathstudent55):

A Pythagorean triple is a set of three numbers that satisfies the Pythagorean equation: \(a^2 + b^2 = c^2 \)

OpenStudy (anonymous):

@mathstudent55 thank you

OpenStudy (mathstudent55):

wlcm